Part A - The vapor pressure of water at 25∘C and 1 atm is about 18 torr. What is the approximate vapor pressure of water at 80∘ ∘ C?
Answer #1

Given,

Vapour pressure of water at 25 oC (P1) = 18 Torr

Vapour pressure of water at 80oC (P2) = ?

Temperature(T1) = 25 oC + 273.15 = 298.15 K

Temperature(T2) = 80 oC + 273.15 = 353.15 K

Now, We know, the Clausius-Claperyon equation,

ln[P1 /P2] = Hvap / R [ 1/T2 - 1/T1]

Here, Hvap of water = 40700 J/mol

R = 8.314 J/K.mol

Substituting the known values in the above formula,

ln[18 /P2] = 40700 J/mol / 8.314 J/K.mol [ 1/353.15 K - 1/ 298.15 K]

ln[18 /P2] = 4895.357 x -0.00052

[18 /P2] = e-2.55713

P2 = 232.2 Torr

Thus, the vapour pressure of water at 80 oC is 230 Torr [ 2 s.f]